Q1) According to a Pew Research Center study, in May 2011, 32% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 328 community college students at random and finds that 132 of them have a smart phone. Then in testing the hypotheses:

H0: p = 0.32 versus

Ha: p > 0.32,

what is the test statistic?

z = ________. (Please round your answer to two decimal places.)
Answer #1

phat = 132/328 = 0.4024

n = 328

z = (phat - p)/sqrt(p*(1-p)/n)
= (0.4024 - 0.32)/sqrt(0.32*(1-0.32)/328)
= 3.20